OpenStudy (anonymous):

Solve: x^2 = 21. If there is more than one answer, write your answers separated by a comma. If necessary, round your answer to the nearest tenth.

OpenStudy (jim766):

take the square root of both sides

OpenStudy (jim766):

what is the sq rt of x^2

OpenStudy (anonymous):

I have to write it into like an essay and I have no idea what to do

OpenStudy (jim766):

first just figure out how to solve it. take the square root of both sides what is the sq rt of x^ what is the sq rt of 21

OpenStudy (jim766):

x^2

OpenStudy (jim766):

what times what = x ^2

OpenStudy (anonymous):

square root of 21 is 4. 58

OpenStudy (jim766):

right and the square root of x^2 = x x^2 = 21 x = 4.6 it says to round to the nearest thenth

OpenStudy (anonymous):

Oh okay so the anwer is 4.6 ?

OpenStudy (jim766):

it will also be - 4.6

OpenStudy (jim766):

does that make sense?

OpenStudy (anonymous):

Yes, thanks ! :)